# The Role of Inclusion, Control, and Ownership in Workplace AI-Mediated Communication

### 13 Dec 2023 v2

@@ 4.2. Control
Figure 3 illustrates the frequency of different survey responses for each gender (different columns) for the following questions: how much control did you feel over the process of writing the message (“Control During Process” on the left), and how much control did you feel over the final version of the message (“Control Over Message” on the right). Figure 3a shows that in the condition with the hesitant AI writing style suggestions (left half), during the writing process, 61.1% of the men (first column) and 61.2% of WNBSD participants (second column) reported feeling a lot or great deal of control (the darker blue and darkest blue bars in the columns). In the condition with the self-assured AI writing style (right half), 48% of men and 49% of the WNBSD participants expressed a lot or a great deal of control during the writing process. A two-way ANOVA revealed that the difference between the AI treatment groups was significant (p=0.0044), meaning that participants in the hesitant condition reported more control over the process. There were no significant differences for gender or the interaction between gender and the AI treatment groups for this question. In both treatment conditions, participants generally reported higher levels of control over the final version of the message than the writing process.
In the condition with the hesitant AI writing style (Figure 3b, left), 73% of men and 80.2% of WNBSD participants reported a lot or a great deal of control over the final version of the message. Figure 3b (right) that shows 59.6% of men and 70.7% of WNBSD participants who wrote with a self-assured writing style model reported a lot or a great deal of control over the final version of the message. For this question as well, a two-way ANOVA revealed that the difference between the AI treatment groups was significant (p=0.046), with participants in the hesitant AI writing style condition reporting greater control over the final version of the message. At the same time, there were no significant differences in relation to gender or interactions between gender and the AI treatment group for this question.
@@ 4.3. Ownership
Figure 4 illustrates the frequency of different survey responses for each gender condition (columns) for the following questions: to what extent do you feel like the message you wrote is yours (“Message Is Yours”), and thinking back on the message writing activity, how much did the message sound like you (“Message Sounds Like You”). In the hesitant AI writing style condition (Figure 4a, left), 55.6% of the men and 53.5% of WNBSD participants expressed a lot or a great deal of ownership over the message (the dark blue and darkest blue bars in the column). In the self-assured AI writing style condition (Figure 4a, right), participants experienced a higher level of ownership, with 41.4% of men and 44.1% of the WNBSD participants expressing a similar degree of ownership over the message. A two-way ANOVA revealed that this difference between the AI treatment groups was significant (p=). There was no significant difference for gender or the interaction between gender and the AI treatment groups for this question.
@@ 4.4. Exploratory Analysis: Inclusion and Agency
Figure 4(b) explores the interaction between inclusion (x-axis) and gender. The graph compares men (blue line) and minoritized genders (WNBSD, light blue line). While inclusion and agency have a direct relationship for all gender identities, the slope for WNBSD participants is greater than the slope for men, indicating that perceived inclusion is more strongly tied to perceived agency for minoritized genders.
@@ 5. Discussion
By introducing the construct of inclusion, our work extends the previous research on AI-MC, which has examined perceptions of agency while interacting with an AI writing assistant (Lee et al., 2022; Mieczkowski, 2022; Mieczkowski et al., 2021). We measure the concept of inclusion after the co-writing process by asking participants if they could envision themselves in the model and whether the model’s writing style matched their own. We show that this concept contributes to the perception of agency, a combined metric reflecting ownership and control over the writing process and the final product.
@@
Informed by these findings, Figure 6 presents a new conceptual model of the factors contributing to agency in AI co-writing environments. This model is tentative: not all of these concepts were measured, let alone tested, in our work here. The model suggests that there are two paths for feeling agency in the co-writing process and its outcome. On the top, AI writing style affects AI-user alignment—which we define as the similarity between the AI writing style and the user’s own preferences that may be derived from their identity. High AI-user alignment leads to increased perceived inclusion, positively contributing to perceived agency. In our research, we did not measure AI-user alignment. Instead, we built on the literature suggesting that minoritized genders are less likely to be aligned with assertive AI suggestions – which was not the case in this study. At the bottom of the model, AI writing style affects AI-task alignment, which we define as the perceived usefulness of the AI assistant for the task (e.g., writing to ask for a promotion with an assertive AI). If there is high AI-task alignment, the user will rely on the model to complete the task. However, such reliance, as we have shown, can negatively contribute to the user’s perceived agency. The conceptual model includes an interaction between AI reliance and inclusion, as suggested by our exploratory results.
@@
Our model raises several open questions that have implications for designers. The most salient question is the potential competition between AI-User alignment and AI-Task alignment. Or, how can developers strike a balance between models that are inclusive of various users but are also effective in completing the task? We note that even if the models align with both the user and task, there is some conflict: the AI-User alignment will ultimately contribute to agency, but the AI alignment with the task, which will increase reliance, will reduce agency. This finding complicates the question of how developers can build models that help the user accomplish the task while also supporting the user’s agency when there is tension between those. Further work is needed to understand the trade-off between agency and task performance in real-world applications.
@@
The idea of inclusion via AI-User alignment captures our hypothesis that gender identity will play a role in perceived inclusion and agency. We hypothesized that minoritized genders would feel more included in the hesitant AI writing style condition, at least compared to men. However, our main findings did not show such gender effects. Although existing literature highlights language use differs by gender in negotiation contexts, it often relies on experimental methods (such as laboratory studies) that encompass factors beyond the scope of our experiment like job status, vocal pitch, or physical presence (Amanatullah and Morris, 2010; Mazei et al., 2015). Many of the laboratory studies show it is the small details, like the tone of voice or the wording, which makes gender differences salient (Thimm et al., 2008). The absence of physical cues, as shown in computer-mediated negotiations, reduces gender-based communication bias (Mazei et al., 2015).
@@
Our work explored the impact of AI auto-complete suggestions on feelings of inclusion, control, and ownership over the written text and the implications for users of various backgrounds. The work has a number of limitations. First, in asking people to write as if asking for a promotion, our study focused on an important and realistic workplace task. However, other workplace writing tasks should also be studied to validate the results. Furthermore, we did not measure individual differences in assertiveness, which could have provided additional insight on the effect of the treatment. Pre-task measurement of such attitudes, though, could have introduced bias to the experimental task. We focus on English-speaking US residents. Additional work is needed to understand how users from various linguistic backgrounds and cultures interact with such AI agents. Despite these limitations, this short paper showed some of the first evidence that minoritized genders’ perceptions of control and ownership could vary in some ways from men’s as a response to writer-style biases. We offer a new conceptual model for understanding the impact of these factors on people’s feelings of overall agency in AI-mediated environments.
@@ A.1. Validating the Treatment
We used human annotators and computational methods to validate the effect of the auto-complete suggestions. We are able to analyze the auto-complete suggestions as well as the final message, as our writing app logs all auto-complete suggestions seen by users. For our analysis, we randomly sampled 100 participants from each AI treatment condition. We selected suggestions that appeared at the beginning, middle, and end of the writing task for a total of five suggestions per participant. We used LIWC-22 on the auto-complete suggestions to analyze the following features: positive tone, negative tone, pro-social behavior, politeness, conflict, clout, authenticity, and social behavior as a whole. Clout refers to the social status or confidence people display in writing. Authenticity refers to how much a person self-regulates, with low authenticity indicating the person is socially cautious and high authenticity indicating a person is spontaneous. Based on prior literature (Mullany, [n. d.]; Merchant, 2012; Thimm et al., 2008), self-assured language and hesitant language should differ along these metrics. With the exception of the positive tone metric, t-tests revealed significant differences between the means across all other the metrics (negative tone: t=-6.73, p=; pro-social behavior: t=3.29, p=0.0012; politeness: t= 3.77, p=0.0024, conflict: t=-2.27, p=0.025; clout: t=3.15, p=0.002; authenticity: t=-4.00, p=; social behavior: t=3.24, p=0.001). Table 2 shows several of the suggestions with some selected self-assured suggestions highlighted in dark blue and selected hesitant suggestions highlighted in light blue.
@@
Additional analysis showed the suggestions had an effect on the final message. We had three human annotators evaluate all of the final messages on a scale from 1-10, where 1 indicated no assertiveness and 10 indicated high assertiveness. T-tests revealed the means between conditions to be significant (self-assured/control: t=14.56, p=; self-assured/hesitant: t= 14.56, p=; control/hesitant: t= 10.09, p=). For the LIWC-22 analysis, t-tests showed significant differences between groups, with the exception of the positive tone metric and politeness (negative tone: t=-7.09, p=; pro-social behavior: t=4.83, p=; conflict: t=-2.24, p=0.026; clout: t= 3.00, p=0.003; authenticity: t= -2.24, p=0.03; social behavior: t= 5.36, p=).

### 26 Feb 2024 v3

@@ 4.2. Control
Figure 3 shows the frequency of different survey responses for each gender (different columns) for the following questions: how much control did you feel over the process of writing the message (“Control During Process” on the left), and how much control did you feel over the final version of the message (“Control Over Message” on the right). In the condition with the hesitant AI writing style suggestions (left side of Figure 3a), 61.1% of the men (first column) and 61.2% of WNBSD participants (second column) reported feeling a lot or great deal of control (the darker blue and darkest blue bars in the columns) during the writing process. In the condition with the self-assured AI writing style (right side of Figure 3a), 48% of men and 49% of the WNBSD participants expressed a lot or a great deal of control during the writing process. A two-way ANOVA revealed that the difference between the AI treatment groups was significant (p=0.004), meaning that participants in the hesitant condition reported more control over the process. There were no significant differences for gender or the interaction between gender and the AI treatment groups for this question.
Looking at Figure 3b and control over the final version, in the condition with the hesitant AI writing style (left), 73% of men and 80.2% of WNBSD participants reported a lot or a great deal of control over the final version of the message. The right side of Figure 3b shows 59.6% of men and 70.7% of WNBSD participants who wrote with a self-assured writing style model reported a lot or a great deal of control over the final version of the message. For this question as well, a two-way ANOVA revealed that the difference between the AI treatment groups was significant (p=0.046), with participants in the hesitant AI writing style condition reporting greater control over the final version of the message. At the same time, there were no significant differences in relation to gender or interactions between gender and the AI treatment group for this question. In both treatment conditions, participants generally reported higher levels of control over the final version of the message (Figure 3b) than the during the writing process (Figure 3a).
@@ 4.3. Ownership
Figure 4 illustrates the frequency of different survey responses for each gender condition (columns) for the following questions: to what extent do you feel like the message you wrote is yours (“Message Is Yours”), and thinking back on the message writing activity, how much did the message sound like you (“Message Sounds Like You”). In the hesitant AI writing style condition (left side of Figure 4a), 55.6% of the men and 53.5% of WNBSD participants expressed a lot or a great deal of ownership over the message (the dark blue and darkest blue bars in the column). In the self-assured AI writing style condition (right side of Figure 4a), participants experienced a higher level of ownership, with 41.4% of men and 44.1% of the WNBSD participants expressing a similar degree of ownership over the message. A two-way ANOVA revealed that this difference between the AI treatment groups was significant (p<0.001). There was no significant difference for gender or the interaction between gender and the AI treatment groups for this question.
@@ 4.4. Exploratory Analysis: Inclusion and Agency
Figure 4(b) explores how the interaction between inclusion (x-axis) and gender impacts agency. The graph compares men (blue line) and minoritized genders (WNBSD, light blue line). While inclusion and agency have a direct relationship for all gender identities, the slope for WNBSD participants is greater than the slope for men, indicating that perceived inclusion is more strongly tied to perceived agency for minoritized genders.
@@ 5. Discussion
By introducing the construct of inclusion, our work extends the previous research on AI-MC, which has examined perceptions of agency while interacting with an AI writing assistant (Mieczkowski, 2022; Mieczkowski et al., 2021). We measure the concept of inclusion after the co-writing process by asking participants if they could envision themselves in the model and whether the model’s writing style matched their own. We show that this concept contributes to the perception of agency, a combined metric reflecting ownership and control over the writing process and the final product.
@@
Informed by these findings, Figure 6 presents a new conceptual model of the factors contributing to agency in AI co-writing environments. This model is tentative: not all of these concepts were measured, let alone tested, in our work here. The model suggests that there are two paths for feeling agency in the co-writing process and its outcome. On the top, the AI writing style affects AI-user alignment—which we define as the similarity between the AI writing style and the user’s own preferences that may be derived from their identity. High AI-user alignment leads to increased perceived inclusion, positively contributing to perceived agency. In our research, we did not measure AI-user alignment. Instead, we built on the literature suggesting that minoritized genders are less likely to be aligned with assertive AI suggestions (Merchant, 2012; Herring and Martinson, 2004)—which was not the case in this study. At the bottom of the model, the AI writing style affects AI-task alignment, which we define as the perceived usefulness of the AI assistant for the task (e.g., writing to ask for a promotion with an assertive AI). If there is high AI-task alignment, the user will rely on the model to complete the task. However, such reliance, as we have shown, can negatively contribute to the user’s perceived agency. The conceptual model includes an interaction between AI reliance and inclusion, as suggested by our exploratory results.
@@
Our model raises several open questions that have implications for designers. The most salient question is the potential competition between AI-User alignment and AI-Task alignment. Or, how can developers strike a balance between models that are inclusive of various users but are also effective in completing the task? We note that even if the models align with both the user and task, there is remaining conflict: the AI-User alignment will ultimately contribute to agency, but the AI alignment with the task—which will increase reliance—will also reduce agency. This finding complicates the question of how developers can build models that help the user accomplish the task while also supporting the user’s agency, given the tension between those. Further work is needed to understand the trade-off between agency and task performance in real-world applications.
@@
The idea of inclusion via AI-User alignment captures our hypothesis that gender identity will play a role in perceived inclusion and agency. We hypothesized that minoritized genders would feel more included in the hesitant AI writing style condition, at least compared to men. However, our main findings did not show such gender effects. Although existing literature highlights language use differs by gender in negotiation contexts, it often relies on experimental methods (such as laboratory studies) that encompass factors beyond the scope of our experiment like job status, vocal pitch, or physical presence (Amanatullah and Morris, 2010; Mazei et al., 2015). Many of the laboratory studies show it is the small details, like the tone of voice or the wording, which makes gender differences salient (Thimm et al., 2003). The absence of physical cues, as shown in computer-mediated negotiations, reduces gender-based communication bias (Stuhlmacher and Walters, 1999).
@@
Our work explored the impact of AI auto-complete suggestions on feelings of inclusion, control, and ownership over the written text and considered the implications for users of various backgrounds. The work has a number of limitations. First, in asking people to write as if asking for a promotion, our study focused on an important and realistic workplace task. However, other workplace writing tasks should also be studied to generalize the results. Furthermore, we did not measure individual differences in assertiveness, which could have provided additional insight on the effect of the treatment. Pre-task measurement of such attitudes, though, could have introduced bias to the experimental task. We focus on English-speaking US residents. Additional work is needed to understand how users from various linguistic backgrounds and cultures interact with such AI agents. Despite these limitations, this short paper showed some of the first evidence that minoritized genders’ perceptions of control and ownership could vary in some ways from men’s as a response to writer-style biases. We offer a new conceptual model for understanding the impact of these factors on people’s feelings of overall agency in AI-mediated environments.
Acknowledgements.
This material is based upon work supported by the National Science Foundation under Grant No. CHS 1901151/1901329.References
- (1)
- Amanatullah and Morris (2010) Emily Amanatullah and Michael Morris. 2010. Negotiating Gender Roles: Gender Differences in Assertive Negotiating Are Mediated by Women’s Fear of Backlash and Attenuated When Negotiating on Behalf of Others. Journal of personality and social psychology 98 (2010), 256–67. https://doi.org/10.1037/a0017094
- Amanatullah and Tinsley (2013) Emily T. Amanatullah and Catherine H. Tinsley. 2013. Ask and Ye Shall Receive? How Gender and Status Moderate Negotiation Success. Negotiation and Conflict Management Research 6 (2013), 253–272. https://doi.org/10.1111/ncmr.12014
- Belk (2013) Russell W. Belk. 2013. Extended Self in a Digital World. Journal of Consumer Research 40 (2013), 477–500. https://doi.org/10.1086/671052
- Bender et al. (2021) Emily M. Bender, Timnit Gebru, Angelina McMillan-Major, and Shmargaret Shmitchell. 2021. On the Dangers of Stochastic Parrots: Can Language Models Be Too Big?. In Proceedings of the 2021 ACM Conference on Fairness, Accountability, and Transparency (FAccT ’21). Association for Computing Machinery, New York, NY, USA. https://doi.org/10.1145/3442188.3445922
- Bhat et al. (2023) Advait Bhat, Saaket Agashe, Parth Oberoi, Niharika Mohile, Ravi Jangir, and Anirudha Joshi. 2023. Interacting with Next-Phrase Suggestions: How Suggestion Systems Aid and Influence the Cognitive Processes of Writing. In Proceedings of the 28th International Conference on Intelligent User Interfaces (IUI ’23). Association for Computing Machinery, New York, NY, USA. https://doi.org/10.1145/3581641.3584060
- Bird et al. (2020) Sarah Bird, Miro Dudík, Richard Edgar, Brandon Horn, Roman Lutz, Vanessa Milan, Mehrnoosh Sameki, Hanna Wallach, and Kathleen Walker. 2020. Fairlearn: A toolkit for assessing and improving fairness in AI. Technical Report MSR-TR-2020-32. Microsoft. https://www.microsoft.com/en-us/research/publication/fairlearn-a-toolkit-for-assessing-and-improving-fairness-in-ai/
- Braun et al. (2018) Niclas Braun, Stefan Debener, Nadine Spychala, Edith Bongartz, Peter Sörös, Helge H. O. Müller, and Alexandra Philipsen. 2018. The Senses of Agency and Ownership: A Review. Frontiers in Psychology 9 (2018). https://doi.org/10.3389/fpsyg.2018.00535
- Cesafsky et al. (2019) Laura Cesafsky, Erik Stayton, and Melissa Cefkin. 2019. Calibrating agency: Human-Autonomy Teaming and the future of work amid highly automated systems. In Ethnographic Praxis in Industry Conference Proceedings. Wiley Online Library. https://doi.org/10.1111/1559-8918.2019.01265
- Herring and Martinson (2004) Susan Herring and Anna Martinson. 2004. Assessing Gender Authenticity in Computer-Mediated Language UseEvidence From an Identity Game. Journal of Language and Social Psychology - J LANG SOC PSYCHOL 23 (2004), 424–446. https://doi.org/10.1177/0261927X04269586
- Jakesch et al. (2023) Maurice Jakesch, Advait Bhat, Daniel Buschek, Lior Zalmanson, and Mor Naaman. 2023. Co-Writing with Opinionated Language Models Affects Users’ Views. In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems (CHI ’23). Association for Computing Machinery, New York, NY, USA. https://doi.org/10.1145/3544548.3581196
- Kidd and Birhane (2023) Celeste Kidd and Abeba Birhane. 2023. How AI can distort human beliefs. Science 380 (2023), 1222–1223. https://doi.org/10.1126/science.adi0248
- Koenecke et al. (2020) Allison Koenecke, Andrew Nam, Emily Lake, Joe Nudell, Minnie Quartey, Zion Mengesha, Connor Toups, John R. Rickford, Dan Jurafsky, and Sharad Goel. 2020. Racial disparities in automated speech recognition. Proceedings of the National Academy of Sciences 117 (2020), 7684–7689. https://doi.org/10.1073/pnas.1915768117
- Lee et al. (2022) Mina Lee, Percy Liang, and Qian Yang. 2022. CoAuthor: Designing a Human-AI Collaborative Writing Dataset for Exploring Language Model Capabilities. In Proceedings of the 2022 CHI Conference on Human Factors in Computing Systems (CHI ’22). Association for Computing Machinery, New York, NY, USA. https://doi.org/10.1145/3491102.3502030
- Mazei et al. (2015) Jens Mazei, Joachim Hüffmeier, Philipp Alexander Freund, Alice F Stuhlmacher, Lena Bilke, and Guido Hertel. 2015. A meta-analysis on gender differences in negotiation outcomes and their moderators. Psychological bulletin 141 (2015), 85–104. https://doi.org/10.1037/a0038184
- Mengesha et al. (2021) Zion Mengesha, Courtney Heldreth, Michal Lahav, Juliana Sublewski, and Elyse Tuennerman. 2021. “I don’t Think These Devices are Very Culturally Sensitive.”—Impact of Automated Speech Recognition Errors on African Americans. Frontiers in Artificial Intelligence 4 (2021), 169. https://doi.org/10.3389/frai.2021.725911
- Merchant (2012) Karima Merchant. 2012. How Men And Women Differ: Gender Differences in Communication Styles, Influence Tactics, and Leadership Styles. Claremont McKenna College.
- Mieczkowski et al. (2021) Hannah Mieczkowski, Jeffrey T. Hancock, Mor Naaman, Malte Jung, and Jess Hohenstein. 2021. AI-Mediated Communication: Language Use and Interpersonal Effects in a Referential Communication Task. Proc. ACM Hum.-Comput. Interact. 5, CSCW1, Article 17 (apr 2021), 14 pages. https://doi.org/10.1145/3449091
- Mieczkowski (2022) Hannah N. Mieczkowski. 2022. AI-Mediated Communication: Examining Agency, Ownership, Expertise, and Roles of AI Systems. Ph. D. Dissertation. https://www.proquest.com/dissertations-theses/ai-mediated-communication-examining-agency/docview/2734696274/se-2
- Mirowski et al. (2023) Piotr Mirowski, Kory W. Mathewson, Jaylen Pittman, and Richard Evans. 2023. Co-Writing Screenplays and Theatre Scripts with Language Models: Evaluation by Industry Professionals (CHI ’23). Association for Computing Machinery, New York, NY, USA. https://doi.org/10.1145/3544548.3581225
- Mullany (2007) Louise Mullany. 2007. Gendered discourse in the professional workplace. Springer. https://doi.org/10.1057/9780230592902
- Neff and Nagy (2018) Gina Neff and Peter Nagy. 2018. Agency in the digital age: Using symbiotic agency to explain human-technology interaction. In A Networked Self and Human Augmentics, Artificial Intelligence, Sentience, Zizi Papacharissi (Ed.). Routledge, 11. https://doi.org/10.4324/9781315202082
- Noy and Zhang (2023) Shakked Noy and Whitney Zhang. 2023. Experimental evidence on the productivity effects of generative artificial intelligence. Science 381 (2023), 187–192. https://doi.org/10.1126/science.adh2586
- Poddar et al. (2023) Ritika Poddar, Rashmi Sinha, Mor Naaman, and Maurice Jakesch. 2023. AI Writing Assistants Influence Topic Choice in Self-Presentation. In Extended Abstracts of the 2023 CHI Conference on Human Factors in Computing Systems (CHI EA ’23). Association for Computing Machinery, New York, NY, USA. https://doi.org/10.1145/3544549.3585893
- Ross and Bayer (2021) Morgan Quinn Ross and Joseph B Bayer. 2021. Explicating self-phones: Dimensions and correlates of smartphone self-extension. Mobile Media & Communication 9 (2021), 488–512. https://doi.org/10.1177/2050157920980508
- Ryan and Deci (2017) Richard M Ryan and Edward L Deci. 2017. Self-determination theory: Basic psychological needs in motivation, development, and wellness. Guilford publications.
- Shelby et al. (2023) Renee Shelby, Shalaleh Rismani, Kathryn Henne, AJung Moon, Negar Rostamzadeh, Paul Nicholas, N’Mah Yilla-Akbari, Jess Gallegos, Andrew Smart, Emilio Garcia, and Gurleen Virk. 2023. Sociotechnical Harms of Algorithmic Systems: Scoping a Taxonomy for Harm Reduction. In Proceedings of the 2023 AAAI/ACM Conference on AI, Ethics, and Society (AIES ’23). Association for Computing Machinery, New York, NY, USA. https://doi.org/10.1145/3600211.3604673
- Singh et al. (2023) Nikhil Singh, Guillermo Bernal, Daria Savchenko, and Elena L. Glassman. 2023. Where to Hide a Stolen Elephant: Leaps in Creative Writing with Multimodal Machine Intelligence. ACM Trans. Comput.-Hum. Interact. 30, Article 68 (sep 2023), 57 pages. https://doi.org/10.1145/3511599
- Stuhlmacher and Walters (1999) Alice F. Stuhlmacher and Amy E. Walters. 1999. Gender differences in negotiation outcome: A meta-analysis. Personnel Psychology 52 (1999), 653–677. https://www.proquest.com/scholarly-journals/gender-differences-negotiation-outcome-meta/docview/220142492/se-2
- Thimm et al. (2003) Caja Thimm, Sabine C Koch, and Sabine Schey. 2003. Communicating gendered professional identity: Competence, cooperation, and conflict in the workplace. The handbook of language and gender (2003), 528–549. https://doi.org/10.1002/9780470756942.ch23
- Wan et al. (2023) Yixin Wan, George Pu, Jiao Sun, Aparna Garimella, Kai-Wei Chang, and Nanyun Peng. 2023. “Kelly is a Warm Person, Joseph is a Role Model”: Gender Biases in LLM-Generated Reference Letters. In Findings of the Association for Computational Linguistics: EMNLP 2023. Association for Computational Linguistics. https://doi.org/10.18653/v1/2023.findings-emnlp.243
- Weidinger et al. (2022) Laura Weidinger, Jonathan Uesato, Maribeth Rauh, Conor Griffin, Po-Sen Huang, John Mellor, Amelia Glaese, Myra Cheng, Borja Balle, Atoosa Kasirzadeh, Courtney Biles, Sasha Brown, Zac Kenton, Will Hawkins, Tom Stepleton, Abeba Birhane, Lisa Anne Hendricks, Laura Rimell, William Isaac, Julia Haas, Sean Legassick, Geoffrey Irving, and Iason Gabriel. 2022. Taxonomy of Risks posed by Language Models. In Proceedings of the 2022 ACM Conference on Fairness, Accountability, and Transparency (FAccT ’22). Association for Computing Machinery, New York, NY, USA. https://doi.org/10.1145/3531146.3533088
@@ A.1. Validating the Treatment
We used human annotators and computational methods to validate the effect of the auto-complete suggestions. We are able to analyze the auto-complete suggestions as well as the final message, as our writing app logs all auto-complete suggestions seen by the participants. For our analysis, we randomly sampled 100 participants from each AI treatment condition. We selected suggestions that appeared at the beginning, middle, and end of the writing task for a total of five suggestions per participant. We used LIWC-22 on the auto-complete suggestions to analyze the following features: positive tone, negative tone, pro-social behavior, politeness, conflict, clout, authenticity, and social behavior as a whole. Clout refers to the social status or confidence people display in writing. Authenticity refers to how much a person self-regulates, with low authenticity indicating the person is socially cautious and high authenticity indicating a person is spontaneous. Based on prior literature (Mullany, 2007; Merchant, 2012; Thimm et al., 2003), self-assured language and hesitant language should differ along these metrics. With the exception of the positive tone metric, t-tests revealed significant differences between the means of the self-assured model suggestions and hesitant model suggestions across all the metrics (negative tone: t=-6.73, p<0.001; pro-social behavior: t=3.29, p=0.0012; politeness: t= 3.77, p=0.0024, conflict: t=-2.27, p=0.025; clout: t=3.15, p=0.002; authenticity: t=-4.00, p<0.001; social behavior: t=3.24, p=0.001). Table 2 shows several of the model suggestions with some selected self-assured suggestions highlighted in dark blue and selected hesitant suggestions highlighted in light blue.
@@
Additional analysis showed the suggestions had an effect on the final message. We had three human annotators evaluate all of the final messages on a scale from 1-10, where 1 indicated no assertiveness and 10 indicated high assertiveness. T-tests revealed the means between conditions to be significant (self-assured/control: t=14.56, p<0.001; self-assured/hesitant: t= 14.56, p<0.001; control/hesitant: t= 10.09, p<0.001). Using the same LIWC-22 analysis, t-tests showed significant differences between the text produced by these groups, with the exception of the positive tone metric and politeness (negative tone: t=-7.09, p<0.001; pro-social behavior: t=4.83, p<0.001; conflict: t=-2.24, p=0.026; clout: t= 3.00, p=0.003; authenticity: t= -2.24, p=0.03; social behavior: t= 5.36, p<0.001).